OpenStudy (anonymous):

x^15-64??? Algebra 2 help please

OpenStudy (anonymous):

@kohai

OpenStudy (anonymous):

Simplifying x15 + -64 = 0 Reorder the terms: -64 + x15 = 0 Solving -64 + x15 = 0 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '64' to each side of the equation. -64 + 64 + x15 = 0 + 64 Combine like terms: -64 + 64 = 0 0 + x15 = 0 + 64 x15 = 0 + 64 Combine like terms: 0 + 64 = 64 x15 = 64 Simplifying x15 = 64 Reorder the terms: -64 + x15 = 64 + -64 Combine like terms: 64 + -64 = 0 -64 + x15 = 0

OpenStudy (anonymous):

I am learning about quadratic equations...I know for this one that you would use the equation a^3-b^3=(a-b)(a^2+ab+b^2)

OpenStudy (kohai):

@Britani do you still need help with this one or are you fine?

OpenStudy (anonymous):

i am still stuck

OpenStudy (anonymous):

I don't know what to do with the exponent 15

OpenStudy (kohai):

Using the quadratic formula?

OpenStudy (anonymous):

yeah

OpenStudy (kohai):

Okay, so, what is the quadratic formula?

OpenStudy (anonymous):

For the formula that I am supposed to use is the differences of cubes formula which is a^3-b^3=(a-b)(a^2+ab+b^2)

OpenStudy (kohai):

The difference of cubes formula isn't the same as the quadratic formula.

OpenStudy (anonymous):

I meant the difference of cubes

OpenStudy (anonymous):

@kohai